OpenStudy (anonymous):

A jar contains 6 chocolate chip cookies and 9 peanut butter cookies. Richard grabs 3 cookies at random to pack in his lunch. What is the probability that he drew 2 chocolate chip cookies and 1 peanut butter cookie?

OpenStudy (kirbykirby):

If you are familiar with the hypergeometric distribution, it's basically this. If not, step by step: Since the order of what you pick doesn't matter... then the number of ways to choose 2 chocolate cookies from 6 is \[ {6 \choose 2}\] To get 1 peanut butter cookie, this is \[ {9 \choose 1} \] So, the number of ways to get 2 chocolate and 1 peanut is: \[ {6 \choose 2}{9 \choose 1}\]. Now, the sample space is the total number of cookies you pick out of 15, and there are \[ {15 \choose 3}\] ways to do so. So the probability is: \[ \large \frac{{6 \choose 2}{9 \choose 1}}{{15 \choose 3}}\]

OpenStudy (anonymous):

So, it is \[\frac{ 27 }{ 91 }\]

OpenStudy (kirbykirby):

Another way to think about it: Pretend you are picking Chocolate, CHocolate, Peanut consecutively... 1st draw: P(choco) = 6/15 2nd draw: P(choco) = 5/14, since there is one chocolate missing now 3rd draw: P(peanut) = 9/13, again since another chocolate is missing So P(choco, choco, peanut) = 6/15*5/14*9/13 = 9/91 But this is the arrangement CCP. In fact, we do not care about the order... you chould have also picked: CPC or PCC Since there are 3 more combinations possible, you multiply 9/91 by 3 (You can think of it like P(2 choco and 1 peanut)=Prob(CCP) or Prob(CPC) or Prob(PCC) = 3*Prob(CCP)

OpenStudy (kirbykirby):

yes

OpenStudy (anonymous):

Awesome, thanks. In the first steps, it reminded me of combinations. (6C2*9C1)/15C3

OpenStudy (kirbykirby):

indeed =] yw